OpenStudy (anonymous):

Please help, WILL GIVE MEDAL! The temperature of a gas decreases 0.045°C every minute for 8 min. What is the change in the temperature of the gas? Use a positive or negative number to express the change in the temperature of the gas.

OpenStudy (shae.mccoy):

0.045* every second/ per second 0.045* 8= 9/25= 0.36

OpenStudy (anonymous):

I will medal the first person to answer, please help! :(

OpenStudy (anonymous):

Thanks! So the answer is 0.36?

OpenStudy (shae.mccoy):

Hold on..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Help? Anyone? >_<

OpenStudy (shae.mccoy):

sorry.. Read it wrong

OpenStudy (anonymous):

.36

OpenStudy (anonymous):

?

OpenStudy (bohotness):

iits

OpenStudy (anonymous):

@roz_mar23 ?

OpenStudy (shae.mccoy):

yea.. I was right

OpenStudy (anonymous):

I'm really confused now. >_<

OpenStudy (shae.mccoy):

Every minute about .36c goes down

OpenStudy (shae.mccoy):

its .36

OpenStudy (bohotness):

its

OpenStudy (anonymous):

The answer is .36

OpenStudy (anonymous):

OHHH! I get it...

OpenStudy (bohotness):

-0.36c

OpenStudy (shae.mccoy):

Its not negative

OpenStudy (anonymous):

If it goes down .36 every minute, for 8 minutes, the answer is 2.88. Thanks!

OpenStudy (bohotness):

idk

OpenStudy (shae.mccoy):

Np

OpenStudy (xxcoyotekidxx):

it would not be .36 take a closer look

OpenStudy (anonymous):

Lol, don't know who to medal XD

OpenStudy (anonymous):

@XxCoyoteKidxX ?

OpenStudy (shae.mccoy):

You take a closer look @XxCoyoteKidxX If its not .36 then what is it?

OpenStudy (anonymous):

@Shae.McCoy Yeah?

OpenStudy (xxcoyotekidxx):

it would actually be solved like this .045/8=0.005625 which means it is -0.005625 degrees Celsius per minute

OpenStudy (xxcoyotekidxx):

take a good look at it

OpenStudy (anonymous):

Question closed, @Shae.McCoy is right. Medal for him.

OpenStudy (shae.mccoy):

Im a girl

OpenStudy (anonymous):

Oops, sorry. :P

OpenStudy (bohotness):

cute bunny pic ^

OpenStudy (anonymous):

Thanks, It's one of my favorites! :D

OpenStudy (xxcoyotekidxx):

that can't be correct b/c if it dropped .36 per minute then it couldn't have an amount that it drops that is less than that for a total of 8 minutes look closer @Shae.McCoy

OpenStudy (anonymous):

It dropped .36 in 8 minutes @XxCoyoteKidxX

OpenStudy (xxcoyotekidxx):

wait never mind I read the every minute and the 8 minute backwards.... Sorry guys I've confused myself on this one... word problems sometimes twist my brain around since i got ADHD

OpenStudy (anonymous):

Okay then.... 0_o

OpenStudy (xxcoyotekidxx):

sorry for all the trouble I caused on this one

OpenStudy (anonymous):

It's okay! @XxCoyoteKidxX

OpenStudy (anonymous):

BTW, I have more questions. :P Anyone wanna help?

OpenStudy (anonymous):

@XxCoyoteKidxX

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

Help!

OpenStudy (xxcoyotekidxx):

I'll help

OpenStudy (anonymous):

Four friends are running a marathon relay. The entire race is 26.2 mi, and each friend will run an equal distance. How far will each friend run?

OpenStudy (anonymous):

6.55 mi. each

OpenStudy (xxcoyotekidxx):

each person would run 6.55 miles

OpenStudy (shae.mccoy):

26.2 divided by 4

OpenStudy (shae.mccoy):

6.55

OpenStudy (anonymous):

Thanks! :D I still have one more question.

OpenStudy (xxcoyotekidxx):

go for it

OpenStudy (shae.mccoy):

ok

OpenStudy (anonymous):

I'll post it as a new question so I can give a medal.

OpenStudy (xxcoyotekidxx):

k

OpenStudy (shae.mccoy):

kk

OpenStudy (xxcoyotekidxx):

kkk

OpenStudy (xxcoyotekidxx):

jk